OpenStudy (anonymous):

system of equations (1) x=y-2 4x+y=2 show work thanks (2) y=3x-2 y=2x-5 (3) -4x+5y=17 4x=6y=-6 (4) 8x+3y=4 -7x+5y=-34 (5) 6x+y=-39 3x+2y=-15

OpenStudy (anonymous):

1. x=y-2 4x +y =2 4(y-2) + y= 2 x=(2)-2 4y -8 + y =2 x= 0 5y -8=2 5y=10 y= 2